OpenStudy (anonymous):

115% of what number is 120?

rebeccaxhawaii (rebeccaxhawaii):

add 15 % of 120

OpenStudy (alexandervonhumboldt2):

x*115/100=120 solve for x

OpenStudy (alexandervonhumboldt2):

x=120*100/115=solve

OpenStudy (pawanyadav):

Let 115% of x=120 So 115/100 × x=120 x=120/1.15

OpenStudy (pawanyadav):

=104.3